OpenStudy (anonymous):

Using the equation, Mg (s) + 2HCl (aq) → MgCl2 (aq) + H2 (g), if 24.3 g of Mg and 75.0 g of HCl are allowed to react, calculate the mass of MgCl2 that is produced.

OpenStudy (anonymous):

First you need to calculate the moles of each reactant you have.

OpenStudy (ciarán95):

\[Mg + 2HCl \rightarrow MgCl _{2} + H _{2}\] There is the same number of each type of atom on each side of the equation, so we can verify that it is correctly balances and we can proceed in using it to answer the question. We are told that we used: - 24.3g of Mg - 75g of HCl Here, we only really need to know one of these values in order to answer the question. in either case, we will: - Work out the number of moles of the reactant we have. - Use the balanced chemical equation above to compare the molar ratio between this reactant and the product (in this case MgCl2). - From this, work out the number of moles of the product we will form from the amount of the reactant we used. - Convert this to find the number of grams of the product that is produced. Let's choose Mg as the reactant we will use for this (although we could also equally choose HCl). From the periodic table, the relative atomic mass of Mg is 24.3 g/mole, correct to 1 decimal place (i.e. 1 mole of Mg would have a mass of 24 g). So, to convert 24.3 g to moles we would work out: \[\frac{ mass }{ atomic mass } = \frac{ 24 g }{ 24.3 g/mole } = 1 mole\] Let's now look at the balanced chemical equation, and specifically the numbers in front of the reactants and products (coefficients). The balanced equation tells us that for every 1 mole of Mg we react with every 2 moles of HCl, we will get 1 mole of MgCl2 and 1 mole of H2. So, if we react 1 mole of Mg fully (i.e. there will be none left over and there is sufficient HCl present to react it with, which we assume there is), we should get 1 mole of MgCl2 as a product, meaning that Mg : MgCl2 are in a 1:1 MOLAR RATIO. So, we would expect 1 mole of MgCl2 to form. With the aid of the periodic table once more, the relative molecular mass of MgCl2 is: 24.3 g/mole + 2(35.5 g/mole) = 95.3 g/mole To convert 1 mole of MgCl2 to grams, we work out: \[1 mole \times 95.3 g/mole = 95 g\] which is the mass of the product produced. You can use this stepwise method in soving similar questions. Hope that helps ! :)